# Surrogate models for lightcurves and photosphere properties of Type II supernovae

### v1

@@ 2.2.2 Spectrum surrogate
Although properties such as the bolometric luminosity and photosphere properties like temperature and radius are useful, the bulk of current observations are from surveys with a few bandpasses. We therefore also want to predict photometry from our model. As we discussed above, assuming a blackbody SED is problematic, particularly at late phases when the SN starts to become nebular or at ultra-violet wavelengths where line-blanketing strongly suppresses the flux compared to a blackbody. Here, we find our autoencoder-regressor approach to be sufficiently accurate and efficient. In particular, comparing with a testing set ($25\%$ of all data) kept out of training, we find our autoencoder achieves a reconstruction mean squared error of $0.001$, i.e., we are accurately representing the complex high-dimensional spectrum with just $64$ dimensions. Independently, the regression model achieves a mean squared error of $0.02$ indicating that a feed-forward neural network can accurately learn the mappings between the physical parameters and the latent space. Overall, the loss on the full pipeline is $\approx 5\%$ on the full spectrum.
In Fig. 4, in the top panel we show an example of the predicted photometry from our surrogate model compared to a blackbody SED assumption (assuming the photosphere properties from our surrogates above) and the true photometry from the stella simulations. In the bottom panel, we show the absolute AB magnitude $90\%$ credible interval for the LSST/u and LSST/y bands. The former showcases that the discrepancies between our spectra surrogate model and the true inputs are small and likely well captured by the uncertainties on the true simulations as well as those with our surrogate pipeline (as we will discuss later). Furthermore, the top panel also highlights the large discrepancy between the blackbody SED assumption model and the true simulation photometry, particularly at phases other than from $50-100$ days. Meanwhile our surrogate model provides a strong match to the true photometry from the simulations. The latter provides valuable diagnostics for observers and potentially allow for more optimized follow-up post discovery. We note that our simulated distributions are also broadly consistent with the absolute magnitudes inferred from ZTF in a volume-limited survey (Das et al., 2025), which is promising from the perspective of inferences on real observations.
@@ 3 Validation
As we discussed earlier, we could provide an estimate of the uncertainty from our surrogate model using Monte Carlo sampling, which we could then use directly in our likelihood instead of a constant $\sigma_{\rm sys}$. However, this approach is slower to evaluate, and we find a constant error approach to provide largely identical results. Our simulated data alongside the true input (dashed curves), and the posterior predictions (shaded bands) are all shown in Fig. 8 for different LSST filters. Our posteriors are consistent with both the data and the true input, highlighting the effectiveness of the model to fit the data and correctly recover the input parameters when considering realistic photometric observations. Furthermore, this study also provides us with some expectations of what constraints we can expect from Type II observations in with the Vera Rubin Observatory. In particular, even accounting for our systematic uncertainty on the models themselves, we find that we can recover the $M_{\rm ZAMS}=11.88_{-0.45}^{+0.57}~{}M_{\odot}$, ${}^{56}\rm{Ni}=0.20_{-0.03}^{+0.02}~{}M_{\odot}$, $R_{\rm csm}=3.84_{-1.07}^{+2.34}~{}\times 10^{14}\rm{cm}$ and $\log_{10}\dot{M}=-3.14_{-1.44}^{+0.28}$ $M_{\odot}$yr-1 (all at $68\%$ credible interval on the one-dimensional marginalised posterior), respectively, which are all consistent with the input, i.e., a precision on e.g., the nickel and progenitor ZAMS mass of $\approx 25\%$ and $\approx 9\%$, respectively. We note that observations of Type II SNe in LSST will likely be augmented by follow-up with other telescopes or observed by other surveys and so our constraints could be viewed as a conservative estimate on the expected precision, painting an exciting picture on the expected constraints from Type II SNe in the future.
@@ 4 Application to real observations
Having validated our model on synthetic data with known inputs, we now turn to applying our full photometric model to infer the properties of a few well-observed Type II SNe.
@@ 4.1 SN 2004et
SN 2004et was discovered in September 2004 and classified as a Type II SN (Zwitter et al., 2004), at a distance of $\approx 5.5$Mpc (Li et al., 2005) (although this distance has since been updated). Images from before the SN provided a strong constraint on the progenitor ZAMS mass of $15\pm 2~{}M_{\odot}$ (Li et al., 2005). Although there are also lower progenitor mass estimates of $\approx 8~{}M_{\odot}$ (Crockett et al., 2011). A number of constraints on the progenitor mass and explosion properties have been placed through spectroscopic (Jerkstrand et al., 2012) and hydrodynamical modelling (Sahu et al., 2006; Utrobin & Chugai, 2009; Maguire et al., 2010), with the latter always providing a high-mass estimate of $\geq 20~{}M_{\odot}$ inconsistent with any constraints on the progenitor. Hydrodynamical models also place a constraint on the explosion properties and nickel masses with a 56Ni mass estimate of $0.06\pm 0.04M_{\odot}$ (Maguire et al., 2010) and explosion energies of $1.2\pm 0.3\times 10^{51}$erg.
We model the multi-band photometry up to 250 days [^2] We assumed a distance of 7.72 $\pm$ 0.10 Mpc from Tinyanont et al. (2019) and a reddening of $E(B-V)$ = 0.41 $\pm$ 0.05 mag from Sahu et al. (2006). of SN 2004et from Sahu et al. (2006) with our surrogate model following the same procedure as we do for our synthetic LSST validation. Our lightcurve fits are shown in Fig. 9 which highlights a good fit to the data. Our posteriors on key parameters are shown in Fig. 10. A noteworthy result of our analysis is the consistency of our progenitor mass estimate with the pre-SN images (Li et al., 2005) and the spectral modelling (Jerkstrand et al., 2012). In particular, we infer $M_{\rm ZAMS}=12.15_{-1.06}^{+1.03}M_{\odot}$ which is in contrast to all results based on hydrodynamical models. Meanwhile our constraints on other parameters are all broadly consistent with past results based on hydrodynamical models at $1\sigma$. In particular, we infer a 56Ni mass of $0.09\pm 0.01M_{\odot}$ (Maguire et al., 2010) and explosion energy of $1.7\pm 0.2\times 10^{51}$erg.
@@ 4.2 SN 2012aw
SN 2012aw was discovered in March 2012 (Fagotti et al., 2012) and localised to the M95 galaxy at a distance of $\approx 9.9$Mpc (Bayless et al., 2013; Bose et al., 2013). Again, pre-SN images provide a constraint into the progenitor star with inferred estimates of $12.5\pm 1.5$ (Fraser et al., 2012; Kochanek et al., 2012; Fraser, 2016). Modelling of the lightcurve and spectra provide some constraints on the explosion energy of $1-2\times 10^{51}$erg and a ${}^{56}\rm{Ni}=0.06\pm 0.1$ $M_{\odot}$  (Bose et al., 2013). As with the above case study, we fit all available photometry [^3] We assumed a distance of 9.9 $\pm$ 0.10 Mpc and a reddening of $E(B-V)$ = 0.07 $\pm$ 0.01 mag from Bose et al. (2013). of this SN from Bose et al. (2013) using Redback and our surrogate model, our results are shown in Fig. 9 and 10, which are broadly in agreement with constraints. Our progenitor mass constraint $10.61_{-0.32}^{+0.37}$ is also on the lower end compared to pre-explosion constraints but consistent at $1\sigma$ with the literature (Fraser et al., 2012; Kochanek et al., 2012; Fraser, 2016). However, we do see a departure at late-times from our model posterior prediction on the lightcurve compared to the data, which is perhaps notable as at these epochs the lightcurve should be driven purely by radioactive decay, which is relatively straightforward to model and infer. There is also a discrepancy in our posteriors on the explosion parameters compared to past results. Here, we infer ${}^{56}\rm{Ni}=0.09\pm 0.1$ $M_{\odot}$ and an explosion energy of $0.63_{-0.04}^{+0.05}\times 10^{51}$erg, both inconsistent with the results based on hydrodynamical models (Bose et al., 2013). This could be a product of nebular-phase and spectral features which are missing from the stella simulations.
@@ 4.3 SN 2017gmr
We also analyse SN 2017gmr, a Type II SN discovered in September 2017 and localized to NGC 988 at a luminosity distance of $19.6\pm 1.4$Mpc (Andrews et al., 2019). Notably, SN 2017gmr is much more energetic than typical Type II SNe with estimates from modelling of the lightcurve and spectra suggesting an explosion energy of $\sim 10^{52}$ erg (Utrobin et al., 2021), with a nickel mass of $0.02-0.13M_{\odot}$, higher than either of the above SNe and typical estimates in Type II SNe (0.37$\pm$0.05, Rodríguez et al., 2021, $M_{\odot}$). The early lightcurve also shows hints of CSM interaction (Andrews et al., 2019), which is consistent with systematic analyses of early Type II lightcurves (Förster et al., 2018). Moreover, polarisation measurements strongly hint at an aspherical ejecta profile (Nagao et al., 2019).
We fit the light curves [^4] We assumed a distance of 19.6 $\pm$ 1.4 Mpc and a reddening of $E(B-V)$ = 0.30 $\pm$ 0.05 mag from Andrews et al. (2019). from Andrews et al. (2019) and the results are shown in Fig. 9 and 10. Our results are consistent with the idea that SN 2017gmr is more unique and energetic as we infer posteriors on explosion energy of $3.35_{-0.19}^{+0.18}\times 10^{51}$erg and ${}^{56}\rm{Ni}=0.16\pm 0.01$ $M_{\odot}$, both higher than the previous two SNe we analysed and broadly consistent with past constraints (Utrobin et al., 2021). Our nickel mass constraint here is also higher than constraints from hydro models on a sample of Type II SNe (e.g., Sukhbold et al., 2016; Martinez et al., 2022), but consistent with estimates from other methods (Pejcha & Prieto, 2015). We note that our posterior on $R_{\rm csm}$ strongly rails against the prior at $10^{15}$cm indicating that the model (as currently configured) could provide a better fit and correspondingly shift other inferred parameters.
@@ 5 Conclusions
Focussing more towards the physics, there are significant improvements we could make in the future. Our general framework for building and validating a surrogate model is agnostic to the input simulations themselves and we could build models for other transients or improve the models here for Type II SNe by replacing the input simulations. In particular, we can retrain our models using a different parameterisation of the progenitor parameters to directly connect with stellar evolution. All simulations that were used in this work assumed 56Ni was mixed into $50\%$ of the ejecta, which is likely correlated with the explosion energy (e.g., Eldridge et al., 2019; Kozyreva et al., 2019), which newer models should include. An obvious extension is to also include spectral features such that we can compare our surrogate model directly with high-resolution spectroscopy. Furthermore, at late-times as the supernova becomes nebular, the local-thermodynamic equilibrium assumption built into stella breaks down and instead we could use simulations from software such as sumo (Jerkstrand et al., 2012) or CMFGEN (Hillier & Dessart, 2012), which are more suitable in this regime. We leave these extensions to future work as well as an exploration of our surrogate modelling approach for cosmological applications of Type IIP supernovae.
@@ Acknowledgments
We thank Claes Fransson and Anders Jerkstrand for helpful discussions. N. Sarin and A. Singh acknowledges support from the Knut and Alice Wallenberg Foundation through the “Gravity Meets Light" project. N. Sarin acknowledges the research environment grant “Gravitational Radiation and Electromagnetic Astrophysical Transients” (GREAT) funded by the Swedish Research Council (VR) under Dnr 2016-06012. C. M. B. O. acknowledges support from the Royal Society (grant Nos. DHF-R1-221175 and DHF-ERE-221005). TJM is supported by the Grants-in-Aid for Scientific Research of the Japan Society for the Promotion of Science (JP24K00682, JP24H01824, JP21H04997, JP24H00002, JP24H00027, JP24K00668) and by the Australian Research Council (ARC) through the ARC’s Discovery Projects funding scheme (project DP240101786). Numerical computations were in part carried out on PC cluster at the Center for Computational Astrophysics, National Astronomical Observatory of Japan.

### v2

@@ 2.3.2 Spectrum surrogate
Although properties such as the bolometric luminosity and photosphere properties like temperature and radius are useful, the bulk of current observations are from surveys with a few bandpasses. We therefore also want to predict photometry from our model. As we discussed in Sec. 2.2, assuming a blackbody SED is problematic, particularly at late phases when the SN starts to become nebular or at ultra-violet wavelengths where line-blanketing strongly suppresses the flux compared to a blackbody. Here, we find our autoencoder-regressor approach to be sufficiently accurate and efficient. In particular, comparing with a testing set ($25\%$ of all data) kept out of training, we find our autoencoder achieves a reconstruction mean squared error of $0.001$, i.e., we are accurately representing the complex high-dimensional spectrum with just $64$ dimensions. Independently, the regression model achieves a mean squared error of $0.02$ indicating that a feed-forward neural network can accurately learn the mappings between the physical parameters and the latent space. Overall, the loss on the full pipeline is $\approx 5\%$ on the full spectrum.
In Fig. 4, in the top panel we show an example of the predicted photometry from our surrogate model (dashed curves) compared to a blackbody SED assumption (assuming the photosphere properties from our surrogates above, square curves) and the true photometry from the stella simulations (solid curves). The bottom panel shows $90\%$ credible intervals for LSST u and y bands, demonstrating the range of absolute magnitudes expected for Type II SNe given our trained model. These predictions can enable more optimized follow-up strategies by: (1) predicting optimal observing epochs for maximum signal-to-noise in specific bands, (2) estimating required integration times for spectroscopic follow-up, and (3) helping identify unusual events that deviate from typical Type II behaviour for prioritized observation. We note that our simulated distributions are also broadly consistent with the absolute magnitudes inferred from ZTF in a volume-limited survey (Das et al., 2025), which is promising from the perspective of inferences on real observations. The top panel demonstrates that the discrepancies between our spectra surrogate model and the true inputs are small and likely well captured by the uncertainties on the true simulations as well as those with our surrogate pipeline (as we will discuss later). Furthermore, the discrepancy between the blackbody assumption (square curves) and the true stella simulation highlights that the dangers of assuming a blackbody SED, particularly at phases other than from 5010050-100 days.
@@ 3 Validation
As we discussed earlier, we could provide an estimate of the uncertainty from our surrogate model using Monte Carlo sampling, which we could then use directly in our likelihood instead of a constant $\sigma_{\rm sys}$. However, this approach is slower to evaluate, and we find a constant error approach to provide largely identical results. Our simulated data alongside the true input (dashed curves), and the posterior predictions (shaded bands) are all shown in Fig. 8 for different LSST filters. Our posteriors are consistent with both the data and the true input, highlighting the effectiveness of the model to fit the data and correctly recover the input parameters when considering realistic photometric observations. Furthermore, this study also provides us with some expectations of what constraints we can expect from Type II observations in with the Vera Rubin Observatory. Even accounting for our systematic uncertainty on the model, we recover $M_{\rm ZAMS}=11.88_{-0.45}^{+0.57}~M_{\odot}$, $^{56}\rm{Ni}=0.20_{-0.03}^{+0.02}~M_{\odot}$, $R_{\rm csm}=3.84_{-1.07}^{+2.34}~\times 10^{14}\rm{cm}$ and $\log_{10}\dot{M}=-3.14_{-1.44}^{+0.28}$ $M_{\odot}$yr-1 ($68\%$ credible interval), respectively. These constraints are all consistent with the input and represent strong constraints on key parameters. For example, a precision the nickel and progenitor ZAMS mass of $\approx 25\%$ and $\approx 9\%$, respectively. We note that the simulated observations above used the proposed rubin_baseline_v3.2 observing-log from LSST and represents the lightcurve of a typical Type II SNe in LSST with this survey schedule. As such, the precision above may not be representative for all Type II SNe observed by LSST. However, many Type II SNe observed by LSST will likely have follow-up observations with other telescopes and so our constraints could be viewed as a conservative estimate on the expected precision. This paints an exciting picture on the expected constraints from Type II SNe in the future.
@@ 4 Application to real observations
Having validated our model on synthetic data with known inputs, we now turn to applying our full photometric model to infer the properties of a few well-observed Type II SNe. We selected these three well-observed Type II SNe as they have extensive multi-wavelength observations and have been studied by different groups in the past, allowing us to compare our constraints. These events also probe different parts of the parameter space with regards to explosion energy, which is a critical parameter in the model. Moreover, two of these events have strong constraints on the progenitor mass from pre-explosion imaging, which we can directly compare our constraint to.
@@ 4.1 SN 2004et
SN 2004et was discovered in September 2004 and classified as a Type II SN (Zwitter et al., 2004), at a distance of $\approx 5.5$Mpc (Li et al., 2005) (although this distance has since been updated to 7.72 $\pm$ 0.10 Mpc (Tinyanont et al., 2019)). Pre-explosion imaging from before the discovery provides strong constraints on the progenitor of SN 2004et of $M_{\rm ZAMS}=15\pm 2~M_{\odot}$ (Li et al., 2005). although there are also lower progenitor mass estimates from such observations of $\approx 8~M_{\odot}$ (Crockett et al., 2011). A number of constraints on the progenitor mass and explosion properties have been placed through spectroscopic (Jerkstrand et al., 2012) and hydrodynamical modelling (Sahu et al., 2006; Utrobin & Chugai, 2009; Maguire et al., 2010) of the SN, with the latter always providing a high-mass estimate of $\geq 20~M_{\odot}$ inconsistent with any constraints on the progenitor. Hydrodynamical models also place a constraint on the explosion properties and nickel masses with a 56Ni mass estimate of $0.06\pm 0.04M_{\odot}$ (Maguire et al., 2010) and explosion energies of $1.2\pm 0.3\times 10^{51}$erg.
We model the multi-band photometry up to 250 days [^2] We assumed a distance of 7.72 $\pm$ 0.10 Mpc from Tinyanont et al. (2019) and a reddening of E(BV)E(B-V) = 0.41 $\pm$ 0.05 mag from Sahu et al. (2006). of SN 2004et from Sahu et al. (2006) with our surrogate model following the same procedure as we do for our synthetic LSST validation. Our lightcurve fits are shown in Fig. 9 which highlights a good fit to the data. Our posteriors on key parameters are shown in Fig. 10. A noteworthy result of our analysis is the consistency of our progenitor mass estimate with the pre-SN images (Li et al., 2005) and the spectral modelling (Jerkstrand et al., 2012). In particular, we infer $M_{\rm ZAMS}=12.15_{-1.06}^{+1.03}M_{\odot}$ which is in contrast to all results based on hydrodynamical models. Our constraints on other parameters are all broadly consistent with past results based on hydrodynamical models at $1\sigma$. In particular, we infer a 56Ni mass of $0.09\pm 0.01M_{\odot}$, and explosion energy of $1.7\pm 0.2\times 10^{51}$erg consistent with previous results (Maguire et al., 2010). The consistency of our progenitor mass estimate with pre-explosion imaging but disagreement with other hydrodynamical models could be a product of a few factors. One of the obvious candidates may be the progenitors from Sukhbold et al. (2016), which are different to the progenitors assumed in any of the previous papers. Another candidate may be the inclusion (and parameterisation) of extensive CSM in our grid which is in contrast to many simulations in the past. A larger sample of constraints with our model compared to constraints from hydrodynamical modelling and pre-explosion imaging will likely reveal the exact cause of the discrepancy.
@@ 4.2 SN 2012aw
SN 2012aw was discovered in March 2012 (Fagotti et al., 2012) and localised to the M95 galaxy at a distance of $\approx 9.9$Mpc (Bayless et al., 2013; Bose et al., 2013). Again, pre-SN images provide a constraint into the progenitor star with inferred estimates of $12.5\pm 1.5M_{\odot}$ (Fraser et al., 2012; Kochanek et al., 2012; Fraser, 2016). Modelling of the lightcurve and spectra provide some constraints on the explosion energy of $1-2\times 10^{51}$erg and a $^{56}\rm{Ni}=0.06\pm 0.1$ $M_{\odot}$  (Bose et al., 2013). Similar to the previous case study, we fit all available photometry from Bose et al. (2013) [^3] We assumed a distance of 9.9 $\pm$ 0.10 Mpc and a reddening of E(BV)E(B-V) = 0.07 $\pm$ 0.01 mag from Bose et al. (2013). using Redback and our surrogate model. Our results are shown in Fig. 9 and 10, which are broadly in agreement with constraints. Our progenitor mass constraint $10.61_{-0.32}^{+0.37}M_{\odot}$ is also on the lower end compared to pre-explosion constraints but consistent at $1\sigma$ with the literature (Fraser et al., 2012; Kochanek et al., 2012; Fraser, 2016). However, we do see a departure at late-times from our model posterior prediction on the lightcurve compared to the data, which is perhaps notable as at these epochs the lightcurve should be driven purely by radioactive decay, which is relatively straightforward to model and infer. There is also a discrepancy in our posteriors on the explosion parameters compared to past results. Here, we infer $^{56}\rm{Ni}=0.09\pm 0.1$ $M_{\odot}$ and an explosion energy of $0.63_{-0.04}^{+0.05}\times 10^{51}$erg, both inconsistent with the results based on hydrodynamical models (Bose et al., 2013). The discrepancies likely stem from a combination of a few factors: 1) the late-time departure from our model suggests missing physics in the nebular phase, where the LTE assumption in stella breaks down. 2) Our nickel-mass estimate may be biased by the fixed choice of mixing and (3) the lower explosion energy we infer could reflect differences in how CSM interaction is treated. Future improvements should incorporate non-LTE nebular phase modelling, variable mixing parameters to better match 2012aw-like events, and investigate the role of the CSM with isolated modelling of the early-time data.
@@ 4.3 SN 2017gmr
[…]
We fit the light curves [^4] We assumed a distance of 19.6 $\pm$ 1.4 Mpc and a reddening of E(BV)E(B-V) = 0.30 $\pm$ 0.05 mag from Andrews et al. (2019). from Andrews et al. (2019) and the results are shown in Fig. 9 and 10. Our results are consistent with the idea that SN 2017gmr is more unique and energetic as we infer posteriors on explosion energy of $3.35_{-0.19}^{+0.18}\times 10^{51}$erg and $^{56}\rm{Ni}=0.16\pm 0.01$ $M_{\odot}$, both higher than the previous two SNe we analysed and broadly consistent with past constraints (Utrobin et al., 2021). Our nickel mass constraint here is also higher than constraints from hydro models on a sample of Type II SNe (e.g., Sukhbold et al., 2016; Martinez et al., 2022), but consistent with estimates from other methods (Pejcha & Prieto, 2015). We note that our posterior on $R_{\rm csm}$ strongly rails against the prior at $10^{15}$cm indicating that the model (as currently configured) could provide a better fit (if allowed to extrapolate to higher radii) and correspondingly shift other inferred parameters.
@@ 5 Conclusions
Focussing more towards the physics, there are significant improvements we could make in the future. Our general framework for building and validating a surrogate model is agnostic to the input simulations themselves and we could build models for other transients or improve the models here for Type II SNe by replacing the input simulations. In particular, we can retrain our models using a different parameterisation of the progenitor parameters to directly connect with stellar evolution. All simulations that were used in this work assumed 56Ni was mixed into $50\%$ of the ejecta, this will likely vary in reality. Moreover, the degree of mixing is probably correlated with the explosion energy (e.g., Eldridge et al., 2019; Kozyreva et al., 2019), an effect that newer models should include. An obvious extension is to also include spectral features such that we can compare our surrogate model directly with high-resolution spectroscopy. Furthermore, at late-times as the supernova becomes nebular, the local-thermodynamic equilibrium assumption built into stella breaks down and instead we could use simulations from software such as sumo (Jerkstrand et al., 2012) or CMFGEN (Hillier & Dessart, 2012), which are more suitable in this regime. This limitation likely affects our constraints on SN2012aw, but is valid limitation for any Type II SN with observations at $\gtrsim 100$ days, where these effects become increasingly important.
Despite the current limitations, the general success of the model and the computational advantage of our surrogate model approach will enable efficient population studies of Type II SNe. Moreover, the general approach introduced here can be readily extended to other transient classes, opening new avenues for constraining stellar evolution and explosion physics in the big-data era.
@@ Acknowledgments
We thank Claes Fransson and Anders Jerkstrand for helpful discussions. We also thank the anonymous referee for their thoughtful comments on the manuscript. N. Sarin and A. Singh acknowledges support from the Knut and Alice Wallenberg Foundation through the “Gravity Meets Light" project. N. Sarin acknowledges the research environment grant “Gravitational Radiation and Electromagnetic Astrophysical Transients” (GREAT) funded by the Swedish Research Council (VR) under Dnr 2016-06012. C. M. B. O. acknowledges support from the Royal Society (grant Nos. DHF-R1-221175 and DHF-ERE-221005). TJM is supported by the Grants-in-Aid for Scientific Research of the Japan Society for the Promotion of Science (JP24K00682, JP24H01824, JP21H04997, JP24H00002, JP24H00027, JP24K00668) and by the Australian Research Council (ARC) through the ARC’s Discovery Projects funding scheme (project DP240101786). Numerical computations were in part carried out on PC cluster at the Center for Computational Astrophysics, National Astronomical Observatory of Japan.